### v1

@@ 2.2 Periodic Behavior
Weak proportionality ensures that an apportionment method allocates seats in exact proportion to the votes received when possible. Clearly this occurs when the house size is a multiple of the total number of votes. It may also occur for smaller values of $h$. In Example 2, the smallest number of house seats $h$ for which an exact proportional allocation is possible is $23$ since $p_{1}=32$ and $p_{2}=14$ are both divisible by $2$ and $(p_{1}+p_{2})/2=23.$
More generally, let $(p_{1}+\cdots+p_{n})/\gcd(p_{1},p_{2},\dots,p_{n})$. It follows that the vote vector $\mathbf{P}=\frac{(p_{1},p_{2},\dots,p_{n})}{\gcd(p_{1},p_{2},\dots,p_{n})}$ is an integer vector that is proportional to $\mathbf{p}$. Thus by both homogeneity and weak proportionality, we have the following proposition.
@@
From Example 2, we see that not only is $F_{1}(46,\mathbf{p})=2\cdot F_{1}(23,\mathbf{p})$, but the sequence $S_{1}(46,\mathbf{p})$ consists of the sequence $S_{1}(23,\mathbf{p})$ repeated twice. In fact, this periodicity characterizes stationary divisor methods.
@@
For a fixed $\mathbf{p}$ and cut point $c$, every stationary divisor method corresponds to a periodic sequential apportionment sequence $S_{c}(\mathbf{p})=s_{1},s_{2},\ldots,$ satisfying $s_{mP+k}=s_{k}$ for all $k=1,\ldots,P$ and $m\geq 1$ where $P=(p_{1}+\cdots+p_{n})/\gcd(p_{1},\dots,p_{n})$.
@@
Let $d=\gcd(p_{1},\ldots,p_{n})$ and recall from Proposition 4 that if $\mathbf{P}=\mathbf{p}/d$ then $F_{c}(mP,\mathbf{p})=m\cdot\mathbf{P}$ for all $m\geq 1$. Suppose that $S_{c}(P,\mathbf{p})=(s_{1},\ldots,s_{P})$. We consider which party receives seat $mP+1$. This is awarded to the party maximizing the ratio $\frac{p_{i}}{mP_{i}+c}$ or, equivalently, minimizing the ratios
$\left\{\frac{mP_{i}+c}{p_{i}}\right\}_{i}=\left\{\frac{m}{d}+\frac{c}{p_{i}}\right\}_{i}.$
This is the same party that minimizes $\{\frac{c}{p_{i}}\}_{i}$; hence $s_{mP+1}=s_{1}$.
Seat $mP+2$ seat is awarded to the party that minimizes the ratio
$\left\{\left\{\frac{mP_{i}+c}{p_{i}}\right\}_{i\neq s_{i}},\frac{mP_{s_{1}}+1+c}{p_{s_{1}}}\right\}=\left\{\left\{\frac{m}{d}+\frac{c}{p_{i}}\right\}_{i\neq s_{i}},\frac{m}{d}+\frac{1+c}{p_{s_{1}}}\right\}.$
Again, subtracting the common term $m/d$ this is the same party that received the $2^{nd}$ seat; hence $s_{mP+2}=s_{2}.$
@@
Although we will usually assume that the $p_{i}$ are natural numbers, so that $p_{i}/p_{j}$ is rational, divisor methods can be equally defined for irrational values. One can imagine contexts in which irrational values could arise in the JIT sequencing context. However, when $p_{i}/p_{j}$ is irrational for some $i$ and $j$, the apportionment sequence is not periodic.
@@
Suppose the apportionment sequence is periodic with period $P$ and suppose $F(P,\mathbf{p})=(a_{1},\ldots,a_{n})$ so that $F(mP,\mathbf{p})=(ma_{1},\ldots,ma_{n})$ for all positive integers $m$. Fix $m$ and consider how the $mP+1$ seat would be assigned. If $p_{i}>p_{j}$ then party $i$ receives $ma_{i}+1$ seats before party $j$ receives $ma_{j}+1$ seats, which implies that
@@
Thus the ratio $p_{i}/p_{j}=a_{i}/a_{j}$ is rational, which is a contradiction. Hence, the apportionment sequence cannot be periodic. ∎
@@
Hill-Huntington’s method corresponds to the rounding rule $f(k)=\sqrt{(k-1)k}$. If $\mathbf{p}=(23,4)$ then the sequences for $h=27$ and $h=54$ are
@@
[…]
Notice that the sequence $S_{H}(54,\mathbf{p})$ is not just the sequence $S_{H}(27,\mathbf{p})$ repeated twice. These two sequences of length 54 agree up to the 28th term, but the 29th term in twice-repeated $S_{H}(27,\mathbf{p})$ is a $2$, while the 29th term in $S_{H}(54,\mathbf{p})$ is a 1. Hence, the sequence is not periodic, despite Hill-Huntington’s method being weakly proportional.
@@ 3 Apportionment Sequences for Two Parties
Suppose that there are two parties with vote totals $p_{1}>p_{2}$. Then Theorem 5 states that the sequential apportionment from any stationary divisor is periodic with length $(p_{1}+p_{2})/\gcd(p_{1},p_{2}).$ The exact sequence depends on the value of $c\in[0,1]$.
To understand what these sequences look like, we assume first that $p_{1}$ and $p_{2}$ are relatively prime, so that $\gcd(p_{1},p_{2})=1$. By weak proportionality and homogeneity, each period of length $p_{1}+p_{2}$ consists of $p_{1}$ 1’s and $p_{2}$ 2s. The following lemma places constraints on what sequences are possible.
@@
For a 2-party apportionment under a stationary divisor method, if $p_{1}>p_{2}$ party 2 never receives two seats in a row.
@@
Fix $0\leq c\leq 1$ and suppose for some $h=a_{1}+a_{2}$, the apportionment is $\mathbf{h}=(a_{1},a_{2})$, and that party 1 receives the next seat followed by 2 seats for party 2. Then
$\frac{p_{1}}{a_{1}-1+c}>\frac{p_{2}}{a_{2}-1+c},\quad\frac{p_{2}}{a_{2}-1+c}>\frac{p_{1}}{a_{1}+c}\quad\mbox{and}\quad\frac{p_{2}}{a_{2}+c}>\frac{p_{1}}{a_{1}+c}.$
The first inequality implies $\frac{(a_{1}-1)p_{2}-(a_{2}-1)p_{1}}{p_{1}-p_{2}}<c$. The third inequality implies $c<\frac{a_{1}p_{2}-a_{2}p_{1}}{p_{1}-p_{2}}$. These two inequalities cannot both be true, so no such $c$ can exist and party 2 cannot get two seats in a row under sequential apportionment. ∎
@@
$S_{c}(p_{1},p_{2})=\underbrace{1,1,\ldots,1}_{k_{1}\text{ times}},2,\underbrace{1,1,\ldots,1}_{k_{2}\text{ times}},2,\ldots,\underbrace{1,1,\ldots,1}_{k_{p_{2}}\text{ times}},2,\underbrace{1,1,\ldots,1}_{p_{1}-\sum_{i}k_{i}\text{ times}}.$
As in Section 2, we use the notation
$S_{c}(p_{1},p_{2})=1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\cdots\,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})},$
@@ 11 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22
Note that the apportionment sequences are constant for $c$ in intervals of the form $\left[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}}\right)$ (the half-open intervals are due to the tie-breaking rule), with a final apportionment sequence in which there is no final 1 corresponding to $c=1$. Additionally, $k_{1},k_{2},\ldots,k_{7}\in\{2,3\}.$ This is due to the fact that $2<\frac{23}{7}<3$; hence for the sequence to remain roughly proportional, party 1 must receive 2 or 3 seats for every seat received by party 2.
As a first step towards generalizing these observations, we identify necessary and sufficient conditions for an apportionment sequence to be the result of a stationary divisor method.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime with $p_{1}>p_{2}$ and the apportionment sequence has the form
$1^{k_{1}}\;2\;1^{k_{2}}\;2\;\ldots\;1^{k_{j}}\;2\;\ldots\;1^{k_{p_{2}}}\;2\;1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
It follows that
@@
    The apportionment sequence corresponds to a stationary divisor method if and only if
@@
    If these conditions are met, then the cut point $c$ lies in the interval
@@
    Moreover, if $C_{c}=\frac{(p_{1}-p_{2})c}{p_{2}}$ then
    $k_{1}$ $=\lfloor C_{c}\rfloor+1,$ (4)
    $k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+C_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+C_{c}\right\rfloor\quad\mbox{ for }i=2,\ldots,p_{2},\text{ and}$
    $k_{p_{2}+1}$ $=p_{1}-\left\lfloor\frac{p_{1}}{p_{2}}(p_{2}-1)+C_{c}\right\rfloor-1.$
We prove Proposition 8 in reverse order.
@@
Suppose the apportionment sequence has the indicated form and that it corresponds to a stationary divisor method with cut point $c$. Then for each $i=1,\ldots,p_{2}$, since party 1 receives $k_{1}+\cdots+k_{i}=\sum_{j\leq i}k_{j}$ seats before party 2 receives $i$ seats, $c$ must satisfy (using the tie-breaking rule),
@@
Thus, $c$ must lie in the interval (3).
To show that the $k_{i}$ satisfy (4), we solve for $\sum_{j\leq i}k_{j}$ in (5) to obtain
@@
Thus, $\sum_{j\leq i}k_{j}=\lfloor\frac{(p_{1}-p_{2})c}{p_{2}}+\frac{p_{1}}{p_{2}}(i-1)\rfloor+1$ for all $i=1,\ldots,p_{2}$. Substituting $i=1$, $i=2$, etc. sequentially into this expression yields the desired result, and proves the second statement.
The proof of the first statement follows. To show that (1) and (2) are necessary, let $i=1$ in (6) to get
$\frac{p_{2}(k_{1}-1)}{p_{1}-p_{2}}\leq c<\frac{p_{2}k_{1}}{p_{1}-p_{2}}.$
Since $c\in[0,1]$ this implies $\frac{p_{2}(k_{1}-1)}{p_{1}-p_{2}}<1$ and $\frac{p_{2}(k_{1})}{p_{1}-p_{2}}>0$. Solving for $k_{1}$ and recalling that $k_{1}$ is an integer, yields $1\leq k_{1}\leq\left\lfloor\frac{p_{1}}{p_{2}}\right\rfloor$.
To show that (2) is necessary, note that in order for $c$ to exist, the intervals in (6) must have nonempty intersection. Thus, each left end-point must be to the left of each right end-point. So for each $i\neq i^{\prime},$
If $i<i^{\prime}$, this implies $p_{1}(i^{\prime}-i)<p_{2}(k_{i+1}+\cdots+k_{i^{\prime}}+1)$. If $i^{\prime}<i$ then $p_{1}(i-i^{\prime})>p_{2}(k_{i^{\prime}+1}+\cdots+k_{i}-1)$. Combining these (and switching the labels $i$ and $i^{\prime}$ in the second inequality so that $i^{\prime}>i$) we get
or
Then (2) follows because the $k_{i}$ are integers.
Conversely, suppose that the apportionment sequence has the given form and that (1) and (2) are satisfied. Then the interval in (3) is non-empty and if $c$ lies in the interval, it clearly satisfies (5) for each $i$. Hence the stationary divisor method with cut point $c$ yields the same sequence. It remains to show that the interval (3) lies in $[0,1].$ To see this, note that for each $i\geq 1$,
Thus, the right-hand side of the interval (3) is greater than 0. Similarly, we can show the left-hand side of the interval (3) is less than or equal to 1, implying that $c\in[0,1]$. ∎
We use Proposition 8 to obtain precise expressions for the exponents as follows.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime and that $p_{1}=ap_{2}+b$ for some integers $a\geq 0$ and $0<b<p_{2}$. If $C_{c}=\frac{(p_{1}-p_{2})c}{p_{2}}=(a-1+\frac{b}{p_{2}})c$ then
$k_{1}$ $=\lfloor C_{c}\rfloor+1$
[…]
$k_{p_{2}+1}$ $=a+b-1+\left\lfloor\frac{b}{p_{2}}(p_{2}-1)+C_{c}\right\rfloor$

### v2

@@ 2.2 Periodic Behavior
In Example 2, the smallest number of house seats $h$ for which there is an exactly proportional allocation is $23$; this follows because $p_{1}=32$ and $p_{2}=14$ are both divisible by $2$ and $(p_{1}+p_{2})/2=23$. From Example 3, not only is the apportionment $F_{1}(46,(32,14))=2\cdot F_{1}(23,(32,14))$, but the sequence $S_{1}(46,(32,14))$ is the sequence $S_{1}(23,(32,14))$ concatenated with itself. Below we see that this happens more generally.
Weak proportionality ensures that an apportionment method allocates seats in exact proportion to the votes received when possible. As noted before, this occurs when the house size is a multiple of the number of votes cast. The smallest house size for which exact proportionality occurs is $P=(p_{1}+\cdots+p_{n})/\gcd(p_{1},p_{2},\dots,p_{n})$. It follows that the vote vector $\mathbf{P}=\frac{(p_{1},p_{2},\dots,p_{n})}{\gcd(p_{1},p_{2},\dots,p_{n})}$ is an integer vector that is proportional to $\mathbf{p}$. Thus, by both homogeneity and weak proportionality, we have the following proposition.
@@
The periodic behavior suggested in Example 3 that $S_{1}(46,(32,14))$ is $S_{1}(23,(32,14))$ repeated twice holds in general. For integer vote vectors $\mathbf{p}$, the following theorem shows that the apportionment sequence is periodic for all stationary divisor methods.
@@
For a fixed $\mathbf{p}$, let $P=(p_{1}+\cdots+p_{n})/\gcd(p_{1},\dots,p_{n})$. For $c\in[0,1]$, the stationary divisor method with cutpoint $c$ generates a periodic apportionment sequence $S_{c}(\mathbf{p})=(s_{1},s_{2},\ldots)$ of period $P$ where $s_{mP+k}=s_{k}$ for all $k=1,\ldots,P$ and $m\geq 1$.
@@
Let $d=\gcd(p_{1},\ldots,p_{n})$ and recall from Proposition 3 that if $\mathbf{P}=\mathbf{p}/d$ then $F_{c}(mP,\mathbf{p})=m\cdot\mathbf{P}$ for all $m\geq 1$. Let $S_{c}(P,\mathbf{p})=(s_{1},\ldots,s_{P})$ be the apportionment sequence for the first $P$ seats. We consider which party receives seat $mP+1$. This is awarded to the minimally indexed party that maximizes the ratio $\frac{p_{i}}{mP+c}$. However, the ratio $\frac{p_{i}}{mP+c}$ is maximized when $\frac{mP+c}{p_{i}}=\frac{m}{d}+\frac{c}{p_{i}}$ is minimized. Because $\frac{m}{d}$ is a constant, then seat $mP+1$ is awarded to the minimally indexed party that minimizes $\frac{c}{p_{i}}$ or maximizes $\frac{p_{i}}{c}$; hence $s_{mP+1}=s_{1}$.
Seat $mP+2$ seat is awarded to the party with the smallest index that minimizes the ratios in the set
$\left\{\left\{\frac{mP+c}{p_{i}}\right\}_{i\neq s_{i}},\frac{mP+1+c}{p_{s_{1}}}\right\}=\left\{\left\{\frac{m}{d}+\frac{c}{p_{i}}\right\}_{i\neq s_{i}},\frac{m}{d}+\frac{1+c}{p_{s_{1}}}\right\}.$
Again, subtracting the common term $m/d$, this is the same party that received the $2^{nd}$ seat; hence $s_{mP+2}=s_{2}$.
@@
Although we will usually assume that the $p_{i}$ are natural numbers, so that $p_{i}/p_{j}$ is rational, divisor methods can be applied for irrational values, too. One can imagine contexts in which irrational values could arise in the JIT sequencing context. However, when $p_{i}/p_{j}$ is irrational for some $i$ and $j$, the apportionment sequence is not periodic.
@@
Let $c$ be the cutpoint for a stationary divisor method. Assume that the apportionment sequence $S_{c}(\mathbf{p})$ is periodic with period $P$ for vote vector $\mathbf{p}$ with $p_{i}/p_{j}$ irrational. Because of the periodicity, if $F_{c}(P,\mathbf{p})=(a_{1},\dots,a_{n})$, then $F_{c}(mP,\mathbf{p})=(ma_{1},\ldots,ma_{n})$ for all positive integers $m$.
Fix $m$ and consider how the $mP+1$ seat would be assigned. If $p_{i}>p_{j}$, then party $i$ receives $ma_{i}+1$ seats before party $j$ receives $ma_{j}+1$ seats, which implies that
@@
It follows that the ratio $p_{i}/p_{j}=a_{i}/a_{j}$ is rational, which is a contradiction. Hence, the apportionment sequence cannot be periodic. ∎
@@
Let $\mathbf{p}=(23,4)$. Hill-Huntington’s method corresponds to the rounding rule $f(k)=\sqrt{(k-1)k}$, where, sequentially, the compared ratios are of the form $p_{i}/\sqrt{a_{i}\left(a_{i}+1\right)}$. Because Hill-Huntington’s method is weakly proportional, the apportionment for $27m$ seats is $(23m,4m)$ for all positive integers $m$. However the first block of $27$ seats is awarded in a different sequence from subsequent blocks of $27$. The apportionment sequence under the Hill-Huntington method for $h=27$ is
@@
Subsequent blocks of $27$ seats match the sequential allocation of the Sainte-Laguë method, which is
$S_{1/2}(27,\mathbf{p})=(1^{3},\,2,\,1^{6},\,2,\,1^{5},\,2,\,1^{6},\,1^{3})\neq S_{H}(27,\mathbf{p}).$
Consequently, $S_{H}(\mathbf{p})$ is $S_{H}(27,\mathbf{p})$ appended with an infinite number of copies of $S_{1/2}(27,\mathbf{p})$. This follows because as $h$ increases, the Hill-Huntington method approaches the Sainte-Laguë method. To see this, we rewrite $\sqrt{x(x+1)}$ as $x\sqrt{1+\frac{1}{x}}$ and use the Taylor series to show that the difference between the arithmetic mean $\frac{x+(x+1)}{2}$ and the geometric mean $\sqrt{x(x+1)}=x\sqrt{1+\frac{1}{x}}$ goes to zero as $x$ increases; specifically,
$\lim_{x\rightarrow\infty}x+\frac{1}{2}-\sqrt{x(x+1)}$ $=\lim_{x\rightarrow\infty}x+\frac{1}{2}-x\left(1+\frac{1}{2x}-\frac{1}{8x^{2}}+O\left(x^{-3}\right)\right)$
$=\lim_{x\rightarrow\infty}\frac{1}{8x}-O\left(x^{-2}\right)=0.$
@@ 3 Apportionment Sequences for Two Parties
For two parties with vote totals $p_{1}>p_{2}$, Theorem 2 states that the apportionment sequence from any stationary divisor is periodic with length $(p_{1}+p_{2})/\gcd(p_{1},p_{2})$. The exact sequence depends on the value of $c\in[0,1]$.
To understand what these sequences look like, we assume first that $p_{1}$ and $p_{2}$ are relatively prime, so that $\gcd(p_{1},p_{2})=1$. By weak proportionality and homogeneity, the apportionment sequence consists of blocks of period length $p_{1}+p_{2}$ and each block consists of $p_{1}$ 1s and $p_{2}$ 2s. The following lemma places constraints on what sequences are possible.
@@
For a 2-party apportionment sequence under a stationary divisor method, if $p_{1}>p_{2}$, then party 2 never receives two seats in a row.
@@
Fix $0\leq c\leq 1$. For the stationary divisor method with cutpoint $c$ and a house size of $h$, assume the apportionment is $\mathbf{a}=(a_{1},a_{2})$, so that $h=a_{1}+a_{2}$. Further, assume that party 1 receives the next seat followed by two seats for party 2. By this assumption, the following inequalities must hold:
$\frac{p_{1}}{a_{1}+c}>\frac{p_{2}}{a_{2}+c},\quad\frac{p_{2}}{a_{2}+c}>\frac{p_{1}}{a_{1}+1+c}\quad\mbox{and}\quad\frac{p_{2}}{a_{2}+1+c}>\frac{p_{1}}{a_{1}+1+c}.$
Respectively, the first and third inequalities imply the inequalities
$\frac{a_{1}p_{2}-a_{2}p_{1}}{p_{1}-p_{2}}<c<\frac{a_{1}p_{2}-a_{2}p_{1}-(p_{1}-p_{2})}{p_{1}-p_{2}}.$
Because $p_{1}-p_{2}>0$, the right side of the above inequality is less than the left side, providing a contradiction. Thus, no such $c$ can exist and party 2 cannot get two seats in a row from a stationary divisor method.∎
@@
$S_{c}(p_{1}+p_{2},(p_{1},p_{2}))=\underbrace{1,1,\ldots,1}_{k_{1}\text{ times}},2,\underbrace{1,1,\ldots,1}_{k_{2}\text{ times}},2,\ldots,\underbrace{1,1,\ldots,1}_{k_{p_{2}}\text{ times}},2,\underbrace{1,1,\ldots,1}_{p_{1}-\sum_{i}k_{i}\text{ times}},$
where the last block of 1s of length $p_{1}-\sum_{i}k_{i}$ may be zero. As in Section 2, we use the notation
$S_{c}(p_{1}+p_{2},(p_{1},p_{2}))=1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\ldots\,,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})},$
@@ 11 $1^{2}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22 $1^{2}$ 22 $1^{2}$ 22 $1^{3}$ 22
Note that the apportionment sequences are constant for $c$ in intervals of the form $\left[\frac{\ell}{p_{1}-p_{2}},\frac{\ell+1}{p_{1}-p_{2}}\right)$ (the half-open intervals are due to the tie-breaking rule); there is also an apportionment sequence in which there is no final 1 corresponding to $c=1$. Additionally, $k_{1},k_{2},\ldots,k_{7}\in\{2,3\}$. For the sequence to remain roughly proportional, party 1 must receive 2 or 3 seats for every seat received by party 2 because $2<\frac{23}{7}<3$.
As a first step towards generalizing these observations, we identify necessary and sufficient conditions on the values of $k_{i}$ for an apportionment sequence to be the result of a stationary divisor method.
@@
Let $p_{1}$ and $p_{2}$ be relatively prime with $p_{1}>p_{2}$. Assume an apportionment sequence of length $p_{1}+p_{2}$ has the form
$1^{k_{1}},\,2,\,1^{k_{2}},\,2,\,\ldots\,,1^{k_{j}},\,2,\,\ldots,\,1^{k_{p_{2}}},\,2,\,1^{p_{1}-(k_{1}+\cdots+k_{p_{2}})}.$
@@
    The apportionment sequence is generated by a stationary divisor method if and only if
@@
    If these conditions from inequalities (1) and (2) are met, then the cutpoint $c$ lies in the interval
@@
    Moreover, if $\delta_{c}=\frac{(p_{1}-p_{2})}{p_{2}}c$, then
    $k_{1}$ $=\lfloor\delta_{c}\rfloor+1,$ (4)
    $k_{i}$ $=\left\lfloor\frac{p_{1}}{p_{2}}(i-1)+\delta_{c}\right\rfloor-\left\lfloor\frac{p_{1}}{p_{2}}(i-2)+\delta_{c}\right\rfloor\quad\mbox{ for }i=2,\ldots,p_{2},\text{ and}$
    $k_{p_{2}+1}$ $=p_{1}-\left\lfloor\frac{p_{1}}{p_{2}}(p_{2}-1)+\delta_{c}\right\rfloor-1.$
@@
We prove part 2. The proof of part 1 is in the Appendix. Suppose the apportionment sequence satisfies the conditions in inequalities (1) and (2) and that it is generated by a stationary divisor method with cutpoint $c$. Then for each $i=1,\ldots,p_{2}$, since party 1 receives $k_{1}+\cdots+k_{i}=\sum_{j\leq i}k_{j}$ seats before party 2 receives $i$ seats, $c$ must satisfy (using the tie-breaking rule),
@@
Thus, $c$ must lie in the interval from Equation (3).
To show that the $k_{i}$ satisfy the equations in (4), we solve for $\sum_{j\leq i}k_{j}$ in (5) to obtain
@@
Because the left and right sides of the above inequality differ by 11, then $\sum_{j\leq i}k_{j}=\left\lfloor\frac{(p_{1}-p_{2})c}{p_{2}}+\frac{p_{1}}{p_{2}}(i-1)\right\rfloor+1$ for all $i=1,\ldots,p_{2}$. Substituting $i=1$, $i=2$, etc. sequentially into this expression we obtain the desired result. ∎
We use Proposition 5 to obtain precise expressions for the exponents in the following theorem, whose proof is in the Appendix.
@@
Suppose $p_{1}$ and $p_{2}$ are relatively prime and that $p_{1}=ap_{2}+b$ for some integers $a\geq 1$ and $0<b<p_{2}$. If $\delta_{c}=\frac{(p_{1}-p_{2})}{p_{2}}c=(a-1+\frac{b}{p_{2}})c$, then
$k_{1}$ $=\lfloor\delta_{c}\rfloor+1,$
$k_{i}$ $=\left\{\begin{array}[]{cl}a+1&\mbox{ if }i\in S\\ a&\mbox{ if }i\notin S\end{array}\right.,\mbox{ and}$
$k_{p_{2}+1}$ $=a+b-1-\left\lfloor\frac{b}{p_{2}}(p_{2}-1)+\delta_{c}\right\rfloor$